Two concentric current loops lie in the same plane. The smaller loop has a radius of 3.4cm and a current of 12 A. The bigger loop has a current of20 A . The magnetic field at the center of the loops is found to be zero.
What is the radius of the bigger loop?
Express your answer to two significant figures and include the appropriate units.
Answer #1

Let B1 is magnetic filed due to smaller loop

B2 is the magnetc filed due to bigger loop.

Apply

Bnet = B1 - B2

0 = B1 - B2

B1 = B2

mue*I1/(2*r1) = mue*I2/(2*r2)

I1/r1 = I2/r2

r2 = r1*(I2/I1)

= 3.4*(20/12)

= 5.67 cm <<<<<<<<<<<---------------Answer